Runner A is initially 5.6 km west of a flagpole
and is running with a constant velocity of
8.4 km/h due east. Runner B is initially 4.9
km east of the flagpole and is running with a
constant velocity of 7.5 km/h due west.
How far are the runners from the flagpole
when their paths cross?

Answer by ikleyn(52799)

.
Runner A is initially 5.6 km west of a flagpole and is running with a constant velocity of 8.4 km/h due east.
Runner B is initially 4.9 km east of the flagpole and is running with a constant velocity of 7.5 km/h due west.
How far are the runners from the flagpole when their paths cross?
~~~~~~~~~~~~~~~~~~~~

Lat assume that the flag pole is the zero point on x-axis directed from west to east.


Then the coordinate of the 1st runner in time is  -5.6 + 8.4*t kilometers;

     the coordinate of the 2nd runner in time is   4.9 - 7.5*t kilometers, 

where t is the time from the start, in hours.


The runners are in the crossing point when

    -5.6 + 8.4t = 4.9 - 7.5t.


Simplify this equation and find t

    8.4t + 7.5t = 4.9 + 5.6

        15.9t   =   10.5

            t   =   10.5/15.9 = 0.660377358 of an hour.


The coordinate of the crossing point is  x = -5.6 + 8.4*0.660377358 = -0.052830193 kilometers,  
or 52.83 meters west from the flag pole.   The numbers are approximate.

Solved.